OpenStudy (anonymous):

What is the magnitude and direction of if G(1, -6) and H(8, 3)? magnitude: 11.4 units: direction: 37.87° magnitude: 11.4 units; direction: 52.13° magnitude: 7.62 units; direction 37.87° magnitude: 7.62 units: direction: 52.13°

OpenStudy (lgbasallote):

\[\text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\text{direction} = \tan^{-1} \left(\frac{y_2 - y_1}{x_2 - x_1}\right)\] does that help?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

wonderful